OpenStudy (precal):

calculator question about relative maximum, relative minimum and point of inflection

OpenStudy (precal):

\[f \prime=\cos(lnx)\]

OpenStudy (precal):

I am given that first derivative of a function f'(x)=cos(lnx) I am suppose to identify relative maximum x=4.8104774 I am suppose to find the relative minimum x=.20787958 I am not sure about the point of inflection I took the second derivative and solved for 0 I think I is x=1

zepdrix (zepdrix):

Did you get this for your second derivative?\[\Large\rm f''(x)=-\frac{1}{x}\sin(\ln x)\]

OpenStudy (precal):

I just cheated and put it in my calculator

zepdrix (zepdrix):

lol :)

OpenStudy (precal):

the TI calculator just takes the general command and will create the 2nd and even the third derivative graph of any function.

zepdrix (zepdrix):

That doesn't seem like a good approach -_- hmmm

OpenStudy (precal):

It really does work

OpenStudy (precal):

I just put in the 2nd derivative in the graphing calculator ie y1=1st derivative function given y2=2nd derivative command (it is Math 8) on the graphing calculator I am not sure if I was to locate the zero value and that would be the point of inflection if I have a sign change

zepdrix (zepdrix):

I'm just.. I don't know how to help you if you're just using your calculator :\ It's just weird..

OpenStudy (precal):

That's ok

zepdrix (zepdrix):

x=1 seems correct. Looks like there is another weird point though.. hmmm

zepdrix (zepdrix):

https://www.desmos.com/calculator/s79b4blgj6

zepdrix (zepdrix):

See that other critical point? That sharp dip downward near zero? It's weird too, cause it didn't show up when I did the calculations by hand.. hmm

OpenStudy (precal):

yes and that is the graph I got with my calculator and that is how I located the min and max

OpenStudy (precal):

I am able to request the next derivative using my graphing calculator

zepdrix (zepdrix):

It'll graph the second derivative for you? oo that should help.

zepdrix (zepdrix):

Oh darn, it looks kind of crazy though doesn't it? :P